# A parametric finite element method for a degenerate multi-phase Stefan problem with triple junctions

### v1

@@ 6 Structure-preserving numerical scheme
In this section, we refine the previously proposed fully discrete scheme (4.4) such that the discrete total energy content is preserved at each time step. To this end, we follow the idea by Bao, Garcke, Nürnberg and Zhao [3] which can interpolate between discrete surface clusters at intermediate timestamps. We define $\Gamma_{i}^{h}(t)=\bigcup_{j=1}^{J_{i}}\overline{\sigma_{i,j}^{h}(t)}$ as the interpolating polyhedral surfaces where $\{\sigma_{i,j}^{h}(t)\}_{j=1}^{J_{i}}$ are mutually disjoint $(d-1)$-simplices with vertices $\{\vec{q}_{i,k}^{h}(t)\}_{k=1}^{K_{i}}$ given by
@@
$\vec{\omega}^{m+\frac{1}{2}}_{i},\,\vec{\xi}\right>^{h}_{\Gamma^{m}_{i}}=\left<\vec{\nu}^{m+\frac{1}{2}}_{i},\,\vec{\xi}\right>_{\Gamma^{m}_{i}}\qquad\forall\ \vec{\xi}\in V^{h}(\Gamma^{m}_{i}),\qquad i\in\mathbb{N}_{\leq I_{S}}.$
@@
$\nabla W^{m+1},\,\nabla\varphi\right>_{\Omega}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<\pi^{h}_{i}\left[\frac{\vec{X}^{m+1}_{i}-\vec{\rm id}}{\tau_{m}}\cdot\vec{\omega}^{m+\frac{1}{2}}_{i}\right],\,\varphi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (6.2a)
@@
$\kappa^{m+1}_{\varsigma},\,\xi\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W^{m+1},\,\xi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (6.2b)
@@
$\kappa^{m+1}_{\varsigma}\vec{\omega}^{m+\frac{1}{2}},\,\vec{\eta}\right>^{h}_{\Gamma^{m}}+\left<\varsigma\nabla_{\Gamma}\vec{X}^{m+1},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma^{m}}=0.$ (6.2c)
We observe that the only change from (4.4) to (6.2) is that $\vec{\omega}^{m}$ is now replaced by $\vec{\omega}^{m+\frac{1}{2}}$. Since the latter depends on $\Gamma^{m+1}=\vec{X}^{m+1}(\Gamma^{m})$, the scheme (6.2) leads to a nonlinear system of equations, in contrast to (4.4). We shall show that (6.2) preserves the total energy content at each time step. Before proving this, we recall the following lemma from [3, Lemma 3.1] without its proof (see also [40, Lemma 4.4]).
@@
Let $(W^{m+1},\vec{X}^{m+1},\kappa^{m+1}_{\varsigma})\in S^{m}\times\underline{V}^{h}(\Gamma^{m})\times V^{h}(\Gamma^{m})$ be a solution to (6.2). Then it holds that $v^{m+1}=v^{m}$.
@@
We deduce from Lemma 6.1 that
@@
[…]
@@ 7 Numerical experiments
For the triangulation $\mathscr{T}^{m}$ of the bulk domain $\Omega$, that is used for the bulk finite element space $S^{m}$, we use an adaptive mesh with fine elements close to the interface $\Gamma^{m}$ and coarser elements away from it. The precise strategy is as described in [13, §5.1] and for a domain $\Omega=(-H,H)^{d}$, for some $H>0$, and two integer parameters $N_{c}<N_{f}$ results in elements with maximal diameter approximately equal to $h_{f}=\frac{2H}{N_{f}}$ close to $\Gamma^{m}$, and elements with maximal diameter approximately equal to $h_{c}=\frac{2H}{N_{c}}$ far away from it. Throughout this section, we always set $H=4$. See Figure 3 for an example adaptive bulk mesh in two space dimensions.
@@
We stress that due to the unfitted nature of our finite element approximations, special quadrature rules need to be employed in order to assemble terms that feature both bulk and surface finite element functions. For all the computations presented in this section, we use true integration for these terms, and we refer to [13, 51] for details on the practical implementation. Throughout this section we use (almost) uniform time steps, in that $\tau_{m}=\tau$ for $m=0,\ldots,M-2$ and $\tau_{M-1}=T-t_{m-1}\leq\tau$. In the sequel, we set all the physical parameters to unity, and in particular $\varsigma_{i}=1$ for all $i\in\mathbb{N}_{\leq I_{S}}$, unless otherwise stated.
@@ 7.2 Numerical simulations in 2d
Figure 9: The evolution of half of the maximum extension in $x$-direction of the two phases over time. For the initial radius $0.49$, $0.5$ or $0.51$ (from left to right) of the initially smaller disk. Below we show plots of the evolutions at times $t=0,1,\ldots,T$ with $T=5$, $1$ and $7$, respectively.
@@
Figure 10: (($\beta_{1},\beta_{2},\beta_{3})=(-1,0,1)=$ right bubble, left bubble, outer phase)
@@
Figure 11: (($\beta_{1},\beta_{2},\beta_{3})=(-1,0,1)=$ right bubble, left bubble, outer phase, and with $\varsigma=(1.75,1,1)$). The solution at times $t=0,0.2,0.5,1,2$, and a plot of the discrete energy over time. For this computation, we have $v^{0}=54.6$.
@@
Figure 12: (($\beta_{1},\beta_{2},\beta_{3})=(0.25,0,-0.25)=$ right bubble plus disk, left bubble, outer phase)
@@
Figure 13: (($\beta_{1},\beta_{2},\beta_{3})=(0.25,0,-0.25)=$ right bubble plus disk, left bubble, outer phase)
@@ 7.3 Convergence experiment in 3d
We also perform a convergence experiment for the true solution of (7.4) for $d=3$ over the time interval $[0,0.5]$. For $i=0\to 3$, we set $N_{f}=2^{5+i}$, $N_{c}=4^{i}$ and $\frac{1}{2}K^{0}_{\Gamma}=\hat{K}(i)$, where $(\hat{K}(0),\hat{K}(1),\hat{K}(2),\hat{K}(3))=(770,3074,12290,49154)$, and $\tau=4^{3-i}\times 10^{-3}$. The errors $\|W^{h}-w\|_{L^{\infty}}$ and $\|\Gamma^{h}-\Gamma\|_{L^{\infty}}$ are displayed in Tables 3 and 4.
@@ 7.4 Numerical simulations in 3d
Figure 15: (($\beta_{1},\beta_{2},\beta_{3})=(-0.1,0,0.1)=$ right bubble, left bubble, outer phase)

### v2

@@ 6 Structure-preserving numerical scheme
In this section, we refine the previously proposed fully discrete scheme (4.4) such that the discrete total energy content is preserved at each time step. To this end, we follow the idea by Bao, Garcke, Nürnberg and Zhao [3] which can interpolate between discrete surface clusters at intermediate timestamps. We define $\Gamma_{i}^{h}(t)=\bigcup_{j=1}^{J_{i}}\overline{\sigma_{i,j}^{h}(t)}$ as the interpolating polyhedral surfaces where $\{\sigma_{i,j}^{h}(t)\}_{j=1}^{J_{i}}$ are mutually disjoint (d1)(d-1)-simplices with vertices $\{\vec{q}_{i,k}^{h}(t)\}_{k=1}^{K_{i}}$ given by
@@
$\left<\vec{\omega}^{m+\frac{1}{2}}_{i},\,\vec{\xi}\right>^{h}_{\Gamma^{m}_{i}}=\left<\vec{\nu}^{m+\frac{1}{2}}_{i},\,\vec{\xi}\right>_{\Gamma^{m}_{i}}\qquad\forall\ \vec{\xi}\in V^{h}(\Gamma^{m}_{i}),\qquad i\in\mathbb{N}_{\leq I_{S}}.$
@@
$\left<\nabla W^{m+1},\,\nabla\varphi\right>_{\Omega}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<\pi^{h}_{i}\left[\frac{\vec{X}^{m+1}_{i}-\vec{\rm id}}{\tau_{m}}\cdot\vec{\omega}^{m+\frac{1}{2}}_{i}\right],\,\varphi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (6.2a)
@@
$\left<\kappa^{m+1}_{\varsigma},\,\xi\right>^{h}_{\Gamma^{m}}+\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<W^{m+1},\,\xi\right>^{(h)}_{\Gamma^{m}_{i}}=0.$ (6.2b)
@@
$\left<\kappa^{m+1}_{\varsigma}\vec{\omega}^{m+\frac{1}{2}},\,\vec{\eta}\right>^{h}_{\Gamma^{m}}+\left<\varsigma\nabla_{\Gamma}\vec{X}^{m+1},\,\nabla_{\Gamma}\vec{\eta}\right>_{\Gamma^{m}}=0.$ (6.2c)
We observe that the only change from (4.4) to (6.2) is that $\vec{\omega}^{m}$ is now replaced by $\vec{\omega}^{m+\frac{1}{2}}$. Since the latter depends on $\Gamma^{m+1}=\vec{X}^{m+1}(\Gamma^{m})$, the scheme (6.2) leads to a nonlinear system of equations, in contrast to (4.4). We shall show that, in addition to being unconditionally stable, (6.2) also preserves the total energy content at each time step. Before proving this, we recall the following lemma from [3, Lemma 3.1] without its proof (see also [40, Lemma 4.4]).
@@
$E(\Gamma^{m+1})+\tau_{m}\|\nabla W^{m+1}\|^{2}_{2}\leq E(\Gamma^{m}).$
Moreover, it holds that $v^{m+1}=v^{m}$.
@@
The proof of the stability bound is identical to the proof of Theorem 4.2. In order to prove the energy content preservation, we deduce from Lemma 6.1 that
@@
$=-\tau_{m}\sum_{i=1}^{I_{S}}\left[\beta\right]_{b_{i}^{-}}^{b_{i}^{+}}\left<\pi^{h}\left[\frac{\vec{X}^{m+1}-\vec{\rm id}}{\tau_{m}}\cdot\vec{\omega}^{m+\frac{1}{2}}\right],\,1\right>^{(h)}_{\Gamma^{m}_{i}}=0.$
@@ 7 Numerical experiments
For the triangulation $\mathscr{T}^{m}$ of the bulk domain $\Omega$, that is used for the bulk finite element space $S^{m}$, we use an adaptive mesh with fine elements close to the interface $\Gamma^{m}$ and coarser elements away from it. The precise strategy is as described in [13, §5.1] and for a domain $\Omega=(-H,H)^{d}$, for some $H>0$, and two integer parameters $N_{c}<N_{f}$ results in elements with maximal diameter approximately equal to $h_{f}=\frac{2H}{N_{f}}$ close to $\Gamma^{m}$, and elements with maximal diameter approximately equal to $h_{c}=\frac{2H}{N_{c}}$ far away from it. Throughout this section, we always set $H=4$. See Figure 3 for an example adaptive bulk mesh in two space dimensions. Observe that employing adaptive bulk meshes does not invalidate the stability results in Theorems 4.2 and 6.1.
@@
We stress that due to the unfitted nature of our finite element approximations, special quadrature rules need to be employed in order to assemble terms that feature both bulk and surface finite element functions. For all the computations presented in this section, unless otherwise stated, we use true integration for these terms, and we refer to [13, 51] for details on the practical implementation. Throughout this section we use (almost) uniform time steps, in that $\tau_{m}=\tau$ for $m=0,\ldots,M-2$ and $\tau_{M-1}=T-t_{m-1}\leq\tau$. In the sequel, we set all the physical parameters to unity, and in particular $\varsigma_{i}=1$ for all $i\in\mathbb{N}_{\leq I_{S}}$, unless otherwise stated.
@@ 7.1 Convergence experiment in 2d
As a comparison, we show the results of the same convergence experiment for the schemes (6.2) and (4.4) with mass-lumping in Tables 3 and 4, respectively.
6.2500e-02 1.5398e-01 6.0357e-03 1.0375e-03 2721 256 $10^{-10}$
3.1250e-02 7.7003e-02 2.0175e-03 1.4508e-03 5425 512 $10^{-10}$
1.5625e-02 3.8500e-02 1.5274e-03 1.0748e-03 10869 1024 $10^{-10}$
7.8125e-03 1.9248e-02 8.4669e-04 6.1128e-04 22761 2048 $10^{-10}$
3.9062e-03 9.6238e-03 8.0877e-04 3.9738e-04 50321 4096 $10^{-10}$
Table 3: Convergence test for (7.1) over the time interval $[0,1]$ for the scheme (6.2) with mass-lumping.
6.2500e-02 1.5392e-01 8.8775e-03 3.0074e-03 2713 256 1.3143e-02
3.1250e-02 7.6994e-02 2.7763e-03 4.8992e-04 5433 512 3.4191e-03
1.5625e-02 3.8499e-02 1.4789e-03 8.3211e-04 10877 1024 8.6125e-04
7.8125e-03 1.9248e-02 8.3425e-04 5.5075e-04 22785 2048 2.1483e-04
3.9062e-03 9.6238e-03 8.0541e-04 3.8228e-04 50329 4096 5.3629e-05
Table 4: Convergence test for (7.1) over the time interval $[0,1]$ for the scheme (4.4) with mass-lumping.
@@ 7.2 Numerical simulations in 2d
Figure 9: The evolution of half of the maximum extension in $x$-direction of the two phases over time. For the initial radius $0.49$, $0.5$ or $0.51$ (from left to right) of the initially smaller disk. Below we show plots of the evolutions at times $t=0,1,\ldots,T$ with $T=5$, 11 and $7$, respectively.
@@
Figure 10: $($$\beta_{1},\beta_{2},\beta_{3})=(-1,0,1)=$ right bubble, left bubble, outer phase)
@@
Figure 11: $($$\beta_{1},\beta_{2},\beta_{3})=(-1,0,1)=$ right bubble, left bubble, outer phase, and with $\varsigma=(1.75,1,1)$). The solution at times $t=0,0.2,0.5,1,2$, and a plot of the discrete energy over time. For this computation, we have $v^{0}=54.6$.
@@
Figure 12: $($$\beta_{1},\beta_{2},\beta_{3})=(0.25,0,-0.25)=$ right bubble plus disk, left bubble, outer phase)
@@
Figure 13: $($$\beta_{1},\beta_{2},\beta_{3})=(0.25,0,-0.25)=$ right bubble plus disk, left bubble, outer phase)
@@ 7.3 Convergence experiment in 3d
We also perform a convergence experiment for the true solution of (7.4) for $d=3$ over the time interval $[0,0.5]$. For $i=0\to 3$, we set $N_{f}=2^{5+i}$, $N_{c}=4^{i}$ and $\frac{1}{2}K^{0}_{\Gamma}=\hat{K}(i)$, where $(\hat{K}(0),\hat{K}(1),\hat{K}(2),\hat{K}(3))=(770,3074,12290,49154)$, and $\tau=4^{3-i}\times 10^{-3}$. The errors $\|W^{h}-w\|_{L^{\infty}}$ and $\|\Gamma^{h}-\Gamma\|_{L^{\infty}}$ are displayed in Tables 5 and 6.
@@ 7.4 Numerical simulations in 3d
Figure 15: $($$\beta_{1},\beta_{2},\beta_{3})=(-0.1,0,0.1)=$ right bubble, left bubble, outer phase)